OpenStudy (anonymous):

When –3 is multiplied by the sum of a number and 8, the result is –21. Find the number. A. 3 B. –1 C. 8 D. –2

OpenStudy (anonymous):

3*(n + 8) = 21 Divide both sides by 3: n + 8 = 21/3 = 7 Substract 8 from both sides: n = 7 - 8 = -1 Answer is B